Circular No. 8794

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 A2 (CHRISTENSEN) E. J. Christensen reports the discovery of a comet on CCD images

taken through light cirrus clouds with the 1.5-m reflector in the course of the Mt. Lemmon Survey, noting a moderately condensed 5′′ coma and an 8′′ tail toward p.a. 300^{o} (description from four 30-s co-added images).

2007 UT  α_{2000}  δ_{2000}  Mag.
Jan. 10.41205  9^{h}01^{m}08.^{s}39  +13^{o}28′58′′.1  19.4

The reported astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-A46.

T = 2007 July 8.036 TT  ω = 226.^{o}013
Ω = 304.395 }2000.0
q = 3.06166 AU  i = 40.974

COMET C/2006 P1 (McNAUGHT) Visual total-magnitude and tail-length estimates made in daylight:

Jan. 12.84 UT, −3.9, short (D. W. E. Green, Cambridge, MA, U.S.A., 20×80 binoc.); 12.91, −4.5, 0^{o}.1 (J. Bortle, Stormville, NY, 10×50 binoc.); 12.91, −4.1, ∼ 0^{o}.5 (C. S. Morris, Fillmore, CA, U.S.A., 20×80 binoc.).

SUPERNOVAE 2007D, 2007E, 2007F Three apparent supernovae have been discovered on unfiltered CCD

frames: 2007E by N. Joubert and W. Li (LOSS/KAIT; cf. IAUC 8792); and the others by T. Puckett (2007D with T. Orff, and 2007F with R. Gagliano; cf. IAUC 8792; 0.35-m reflector at Ellijay, GA).

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007D Jan. 9.02 3^{h}18^{m}38.^{s}71 +37^{o}36′26′′.4 18.2 4′′.9 E, 2′′.8 S
2007E Jan. 10.17 2 26 34.41 −15 50 36.5 15.9 15′′.7 E, 14′′.2 N
2007F Jan. 11.40 13 03 15.08 +50 37 07.7 17.3 10′′.1 W, 7′′.4 S

Additional magnitudes for 2007D in UGC 2653: 2005 Dec. 6, [19.4; 2007 Jan. 11.07, 17.9. Additional KAIT magnitudes for 2007E in NGC 921: 2006 Dec. 7.16, [18.2; 2007 Jan. 12.13, 16.0. Additional magnitudes for 2007F in UGC 8162: 2005 Dec. 16, [19.8 (Puckett); 2007 Jan. 12.27, 16.8 (A. Sehgal, Osoyoos, BC, Canada, 0.50-m reflector).

2007 January 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green